problem 2
Date of pricing: November 15, 2019
Time till expiration: four months / Contract expires on March 15, 2020
Current value of an index: 2 803
Continuously compounded interest rate: 4.5 %
Continuously compounded dividend yield: 2.3%
Forward Price of index JKl is 2823.6309

What is the value of the forward contract on January 15, 2020 if:
Forward price of contract with the same underlying assets that expires on March 15, 2020 is 2 790
Forward price of contract with the same underlying assets that expires on May 15, 2020 is 2 980
Answer #1

Sport Rate on 15, January 2020

2803 X {1+(0.045-0.023/6)}= 2813.28

1. VALUE OF FORWARD CONTRACT ON 15 JANUARY, 2020 WHEN FORWARD PRICE IS 2790 ON MARCH 2020.

PRESENT VALUE OF FORWARD PRICE: 2790/{1+(0.045-0.023/6)}= 2779.80

Value of Forward = Spot Rate – PV of Forward price= 2813.28-2779.80=33.48

1. VALUE OF FORWARD CONTRACT ON 15 JANUARY, 2020 WHEN FORWARD PRICE IS 2980 ON MAY 2020.

PRESENT VALUE OF FORWARD PRICE: 2980/{1+(0.045-0.023/3)}= 2958.31

Value of Forward = Spot Rate – PV of Forward price= 2813.28-2958.31= - 145.03